# RS-RAG: Bridging Remote Sensing Imagery and Comprehensive Knowledge with a Multi-Modal Dataset and Retrieval-Augmented Generation Model

### v1

@@ Abstract
Recent progress in Vision-Language Models (VLMs) has demonstrated impressive capabilities across a variety of tasks in the natural image domain. Motivated by these advancements, the remote sensing community has begun to adopt VLMs for remote sensing vision-language tasks, including scene understanding, image captioning, and visual question answering. However, existing remote sensing VLMs typically rely on closed-set scene understanding and focus on generic scene descriptions, yet lack the ability to incorporate external knowledge. This limitation hinders their capacity for semantic reasoning over complex or context-dependent queries that involve domain-specific or world knowledge. To address these challenges, we first introduced a multimodal Remote Sensing World Knowledge (RSWK) dataset, which comprises high-resolution satellite imagery and detailed textual descriptions for 14,141 well-known landmarks from 175 countries, integrating both remote sensing domain knowledge and broader world knowledge. Building upon this dataset, we proposed a novel Remote Sensing Retrieval-Augmented Generation (RS-RAG) framework, which consists of two key components. The Multi-Modal Knowledge Vector Database Construction module encodes remote sensing imagery and associated textual knowledge into a unified vector space. The Knowledge Retrieval and Response Generation module retrieves and re-ranks relevant knowledge based on image and/or text queries, and incorporates the retrieved content into a knowledge-augmented prompt to guide the VLM in producing contextually grounded responses. We validated the effectiveness of our approach on three representative vision-language tasks, including image captioning, image classification, and visual question answering, where RS-RAG significantly outperformed state-of-the-art baselines. By bridging remote sensing imagery and comprehensive knowledge, RS-RAG empowers remote sensing VLMs with enhanced contextual reasoning, enabling them to generate more accurate, informative, and semantically grounded outputs across a wide range of tasks.
Index Terms:
Remote Sensing, Vision Language Model, Retrieval-Augmented Generation, World Knowledge
@@ I Introduction
Remote sensing imagery, as a critical source of information for Earth observation and monitoring, plays an essential role in urban planning [1], agricultural assessment [2], and environmental protection [3]. However, as remote sensing technology advances, the scale and complexity of imagery data have rapidly increased, making it increasingly difficult for traditional manual analysis or image processing methods to meet practical demands. Deep learning methods [4, 5] have significantly improved the accuracy and efficiency of tasks like classification, segmentation and object detection by automatically extracting features from vast amounts of remote sensing data. While these methods have made notable progress, most deep learning models rely predominantly on single-modal visual information, lacking deep semantic understanding of image content. This limitation results in reduced generalization and adaptability, particularly for tasks that require in-depth semantic analysis and comprehensive scene understanding.
@@
The emergence of Vision-Language Models (VLMs) [6, 7, 8, 9, 10, 11] offers a novel solution for the semantic analysis of remote sensing data. By leveraging multimodal fusion techniques, VLMs combine visual features with language information to automatically generate descriptive insights for remote sensing imagery. This semantic enhancement improves image classification and object detection performance while enabling the transformation of recognition results into natural language descriptions, making them more interpretable and accessible for various applications. Additionally, VLMs perform well even in weakly supervised or zero-shot learning scenarios, providing reliable analysis with limited labeled data and thus reducing dependency on extensive data annotations. This cross-modal integration not only enhances the model’s cognitive ability to interpret remote sensing imagery but also facilitates detailed semantic descriptions of complex scenes, paving the way for broader intelligent applications in remote sensing. However, existing remote sensing VLMs primarily focus on identifying image features and providing basic scene descriptions, lacking deeper background understanding of the objects within the images. Particularly when it comes to rich semantic information requiring remote sensing domain expertise or other general world knowledge, such as historical, cultural, and social contexts, these VLM models often struggle to provide comprehensive contextual support.
@@
To address this issue, we first introduce the Remote Sensing World Knowledge (RSWK) dataset, a multimodal remote sensing dataset that contains high-resolution imagery and natural language descriptions for approximately 14,141 well-known locations worldwide from 175 conuties. Unlike most existing remote sensing vision-language datasets that only provide basic descriptions of current scenes, our RSWK dataset will include richer remote sensing domain expertise and world knowledge about the objects within these scenes. For instance, from a remote sensing perspective, it will provide information on surface reflectance, spectral indices, and atmospheric. From a world knowledge perspective, it will include historical background, cultural significance, construction period, and major events. This combination of remote sensing expertise and world knowledge will not only enhance the RSWK dataset’s utility for visual analysis of remote sensing images, but also provide the model with deeper semantic context, overcoming the limitations of traditional datasets and enabling remote sensing VLMs to perform more complex cognitive tasks. Furthermore, our dataset incorporates historical, cultural, and social backgrounds from various countries and regions, allowing VLM models to be trained across diverse geographical and cultural contexts, thereby improving their generalization ability and understanding of different cultural settings.
@@
Figure 1: The construction process of the Remote Sensing World Knowledge (RSWK) dataset begins with collecting landmark data from around the world, followed by extracting geographic information to pinpoint precise location coordinates. Using these coordinates, remote sensing images are acquired, which are further standardized through image processing techniques. Corresponding remote sensing expert knowledge, such as surface temperature, climate, atmospheric conditions, and spectral coefficients, is also included. Additionally, world knowledge is retrieved from online resources, providing detailed background information about the landmarks, including historical context, cultural significance, and major events. This combined information is structured into organized attributes to align image and text data, forming a final multimodal dataset. The resulting RSWK dataset integrates high-resolution images with extensive remote sensing and world knowledge, enabling advanced semantic understanding in remote sensing applications.
@@
Furthermore, to effectively leverage the comprehensive information provided by the RSWK dataset, we propose the Remote Sensing Retrieval-Augmented Generation (RS-RAG) model. RS-RAG is designed to enhance the capacity of vision-language models to generate contextually enriched and knowledge-grounded responses for remote sensing imagery. It operates by integrating external knowledge, both domain-specific and general world knowledge, retrieved from a multimodal knowledge base constructed using the RSWK dataset. The model consists of two main components: (1) the Multi-Modal Knowledge Vector Database Construction module, which encodes satellite imagery and textual descriptions into a shared embedding space using unified image and text encoders to enable efficient cross-modal retrieval; and (2) the Knowledge Retrieval and Response Generation module, which retrieves top-ranked knowledge entries based on image and text queries, re-ranks them via a fused similarity score, and incorporates the selected content into a knowledge-augmented prompt. This prompt is then passed to the vision-language model, enabling it to generate responses that go beyond semantic understanding and reflect deeper background knowledge. By coupling visual input with relevant contextual information, RS-RAG significantly improves the interpretability and accuracy of vision-language outputs, particularly for complex queries involving geospatial, historical, or environmental reasoning. To assess the effectiveness of our proposed model, we construct a lightweight benchmark and conduct comprehensive experiments on three representative tasks: image captioning, image classification, and visual question answering. Results on these tasks demonstrate the model’s ability to produce accurate, context-rich descriptions, deliver semantically informed scene classifications, and provide precise answers to knowledge-intensive queries by leveraging both visual and textual modalities. Through these findings, RS-RAG demonstrates its potential to substantially advance the capabilities of vision-language models in remote sensing, effectively bridging the gap between imagery and comprehensive contextual knowledge. Our main contributions are summarized as follows:
@@
  • We construct the Remote Sensing World Knowledge (RSWK) dataset, a large-scale multimodal benchmark containing 14,141 high-resolution remote sensing images and rich textual descriptions of globally recognized landmarks from 175 countries. The descriptions incorporate both domain-specific knowledge (e.g., land use, temperature, wind direction) and general world knowledge (e.g., historical, cultural, and societal context).
@@
  • We design a lightweight benchmark for evaluating remote sensing vision-language models across three core tasks: image captioning, image classification, and visual question answering. This benchmark enables systematic assessment of both semantic understanding and deeper knowledge-grounded reasoning.
  • Extensive experiments demonstrate that RS-RAG consistently outperforms state-of-the-art vision-language models across all tasks, particularly on queries requiring external world knowledge. These results highlight the effectiveness of RS-RAG in bridging remote sensing imagery with structured knowledge, and point to promising future directions for research on remote sensing vision-language models.
@@ II-A Remote Sensing Multimodal Datasets
Several multimodal datasets [12, 13, 14, 15, 16] have been developed to bridge the gap between vision and language in the remote sensing domain. UCM Captions [17], Sydney Captions [17], and RSICD [18] are among the earliest datasets that provide textual descriptions for remote sensing images. Each image in these datasets is paired with five relatively simple human-written sentences, offering only a basic level of semantic information. To enhance the quality and richness of textual annotations, RSGPT [6] recently introduced RSICap, a high-quality image captioning dataset with detailed human-annotated descriptions of aerial scenes. RSICap serves as a valuable resource for fine-tuning and developing domain-specific vision-language models in remote sensing. Beyond manual annotation, researchers have explored the construction of large-scale datasets using automatic or hybrid approaches applied to existing data sources. For instance, RS5M [19] was created by aggregating 11 publicly available image–text paired datasets along with three large-scale class-level labeled datasets. Captions were generated using BLIP-2, resulting in a diverse and large-scale dataset suitable for training foundational multimodal models. Similarly, RemoteCLIP [20] compiled a large-scale dataset by integrating 10 object detection datasets, 4 semantic segmentation datasets, and 3 remote sensing image–text datasets, enabling contrastive pretraining for cross-modal alignment. In addition, GeoChat introduced the RS Multimodal Instruction Dataset, which incorporates heterogeneous data sources, including three object detection datasets, one scene classification dataset, and a visual question answering dataset focused on flood detection. More recently, FedRSCLIP [21] proposed a new multimodal remote sensing dataset specifically designed for federated learning scenarios, further expanding the applicability of vision-language research in distributed settings. These efforts collectively advance the development of remote sensing VLMs by providing diverse, rich, and large-scale multimodal data resources tailored to various downstream tasks.
@@ II-B Remote Sensing Multimodal Model
With the growing availability of remote sensing multimodal datasets, a number of vision-language models [6, 7, 8, 9, 10, 11] have been proposed to enhance the understanding and interpretation of aerial imagery through natural language. [22] provides the first comprehensive review of vision-language models in remote sensing, systematically summarizing tasks, datasets, and methods, and identifying key challenges and future directions. RSCLIP [23] is a pioneering vision-language model designed for remote sensing scene classification, which leverages contrastive vision-language supervision and incorporates a pseudo-labeling technique along with a curriculum learning strategy to improve zero-shot classification performance through multi-stage fine-tuning. RSGPT [6] represents one of the earliest attempts to build a generative pretrained model for remote sensing. It fine-tunes only the Q-Former and a linear projection layer on the proposed RSICap dataset, achieving notable improvements in both image captioning and visual question answering tasks. Similarly, GeoChat [24] adapts the LLaVA-1.5 architecture and fine-tunes it on its proposed remote sensing multimodal instruction-following dataset, offering multi-task conversational capabilities grounded in high-resolution satellite imagery. In addition, to address the common issue of hallucination in remote sensing vision-language models, a Helpful and Honest Remote Sensing Vision-Language Model [8], named H2RSVLM, is proposed and fine-tuned on the RSSA dataset, the first dataset specifically designed to enhance self-awareness in remote sensing VLMs. More recently, RSMoE [25] was proposed as the first Mixture-of-Experts-based VLM tailored for remote sensing. It features a novel instruction router that dynamically dispatches tasks to multiple lightweight expert LLMs, allowing each expert to specialize in a specific subset of tasks.
Figure 2: Overview of the RSWK dataset. (a) Global distribution of landmarks used in the dataset, with color indicating the number of landmarks per country. (b) Statistical summaries of landmark counts across the top 100 countries (left) and the top 15 most frequent landmark categories (right). (c) A specific example from the RSWK dataset, showcasing the Sydney Opera House, including its satellite imagery, remote sensing domain knowledge, and structured world knowledge.
@@ III Dataset Construction
Most existing remote sensing VLM datasets focus primarily on basic descriptions of objects in a scene, typically providing only information about the objects present. While these datasets serve a purpose in supporting fundamental tasks such as scene classification and image captioning, they fall short in applications requiring complex semantic understanding or contextual awareness. To address this limitation, in our full paper, we would like to propose the Remote Sensing World Knowledge (RSWK) dataset, which encompasses high-resolution remote sensing imagery of well-known locations worldwide, along with domain knowledge and world knowledge described in natural language. This dataset not only fills the gap in knowledge depth and breadth found in current remote sensing datasets but also provides a new foundation for advancing remote sensing technology toward intelligent applications. The core value of remote sensing data lies in its ability to deliver rich geographic and spatial distribution information. By integrating high-resolution imagery from across the globe with detailed domain knowledge and world knowledge, the RSWK dataset extends this value, expanding the application of remote sensing data from traditional foundational tasks to complex scenarios that require deeper semantic understanding.
III-A Data Processing
Fig. 1 illustrates the end-to-end pipeline for constructing the Remote Sensing World Knowledge (RSWK) dataset, which aims to bridge the gap between remote sensing domain knowledge and encyclopedic world knowledge of globally distributed landmarks. This pipeline is designed to automatically curate and align multimodal information from global sources, integrating high-resolution remote sensing imagery, remote sensing expert knowledge, and contextual world knowledge. The process begins with GPT-4o [26], a state-of-the-art large language model, which is employed to generate a comprehensive list of globally recognized landmarks across diverse countries and regions, ensuring broad cultural and geographic coverage that includes both natural and man-made sites of historical and societal relevance. To acquire world knowledge, we utilize the Wikipedia API to extract descriptive information for each landmark, such as historical background, cultural relevance, architectural details, and notable events, with the complete set of fields shown in Table I. Post-processing is performed using DeepSeek [27]to remove low-quality or irrelevant entries. In parallel, the Google Geocoding API is used to obtain precise geographic coordinates for each location; to resolve potential ambiguities in place names, we implement a validation mechanism that cross-references multiple sources and applies regional constraints. High-resolution satellite imagery is retrieved from the ArcGIS Tile Map Service, with spatial resolutions ranging from $0.6$m to $0.15$m, followed by normalization, cropping, and resizing to ensure consistency in image quality and dimensions. Concurrently, remote sensing expert knowledge is derived using Google Earth Engine (GEE) [28], as detailed in Table II. This includes a wide range of satellite-derived geophysical and meteorological variables such as land surface temperature, surface albedo, land cover classification, vegetation indices, and precipitation, which are sourced from authoritative datasets like MODIS, Landsat, and ERA5. By integrating these three complementary modalities, remote sensing imagery, remote sensing expert knowledge, and contextual world knowledge, the RSWK dataset provides a rich foundation for multimodal learning in remote sensing, enabling downstream tasks such as image captioning, image classification, and visual question answering.

### v2

# Remote Sensing Retrieval-Augmented Generation: Bridging Remote Sensing Imagery and Comprehensive Knowledge with a Multi-Modal Dataset and Retrieval-Augmented Generation Model

@@ Abstract
Recent progress in Vision-Language Models (VLMs) has demonstrated impressive capabilities across a variety of tasks in the natural image domain. Motivated by these advancements, the remote sensing community has begun to adopt VLMs for remote sensing vision-language tasks, including image captioning, scene understanding, and visual question answering (VQA). However, existing remote sensing VLMs typically rely on closed-set scene understanding and focus on generic scene descriptions, yet lack the ability to incorporate external knowledge. This limitation hinders their capacity for semantic reasoning over complex or context-dependent queries that involve domain-specific or world knowledge. To address these challenges, we first introduced a multimodal Remote Sensing World Knowledge (RSWK) dataset, which comprises high-resolution satellite imagery and detailed textual descriptions for 14,820 well-known landmarks across 16 categories from 184 countries, integrating both remote sensing domain knowledge and broader world knowledge. Building upon this dataset, we proposed a novel Remote Sensing Retrieval-Augmented Generation (RS-RAG) framework, which consists of two key components. The Multi-Modal Knowledge Vector Database Construction module encodes remote sensing imagery and associated textual knowledge into a unified vector space. The Knowledge Retrieval and Response Generation module retrieves and re-ranks relevant knowledge based on image and/or text queries, and incorporates the retrieved content into a knowledge-augmented prompt to guide the VLM in producing contextually grounded responses. We validate the effectiveness of RS-RAG on our designed benchmark covering three vision-language tasks: image captioning, image classification, and VQA, where RS-RAG significantly outperformed state-of-the-art baselines. By bridging remote sensing imagery and comprehensive knowledge, RS-RAG empowers remote sensing VLMs with enhanced contextual reasoning, enabling them to generate more accurate, informative, and semantically grounded outputs across a wide range of tasks.
@@ I Introduction
Remote sensing imagery, as a critical source of information for Earth observation and monitoring, plays an essential role in urban planning [8], agricultural assessment [23], and environmental protection [33]. However, as remote sensing technology advances, the scale and complexity of imagery data have rapidly increased, making it increasingly difficult for traditional manual analysis or image processing methods to meet practical demands. Deep learning methods [43, 19] have significantly improved the accuracy and efficiency of tasks like classification, segmentation, and object detection by automatically extracting features from vast amounts of remote sensing data. While these methods have made notable progress, most deep learning models rely predominantly on single-modal visual information, lacking deep semantic understanding of image content. This limitation results in reduced generalization and adaptability, particularly for tasks that require in-depth semantic analysis and comprehensive scene understanding.
@@
The emergence of Vision-Language Models (VLMs) [10, 4, 28, 37, 15, 36] offers a novel solution for the semantic analysis of remote sensing data. By leveraging multimodal fusion techniques, VLMs combine visual features with language information to automatically generate descriptive insights for remote sensing imagery. This semantic enhancement improves image classification and object detection performance while enabling the transformation of recognition results into natural language descriptions, making them more interpretable and accessible for various applications. Additionally, VLMs perform well even in weakly supervised or zero-shot learning scenarios, providing reliable analysis with limited labeled data and thus reducing dependency on extensive data annotations. This cross-modal integration not only enhances the model’s cognitive ability to interpret remote sensing imagery but also facilitates detailed semantic descriptions of complex scenes, paving the way for broader intelligent applications in remote sensing. However, existing remote sensing VLMs primarily focus on identifying image features and providing basic scene descriptions, lacking a deeper background understanding of the objects within the images. Particularly when it comes to rich semantic information requiring remote sensing domain expertise or other general world knowledge, such as historical, cultural, and social contexts, these VLMs often struggle to provide comprehensive contextual support.
@@
To address this issue, we first introduce the Remote Sensing World Knowledge (RSWK) dataset, a multimodal benchmark that includes high-resolution imagery and natural language descriptions for over 14,000 well-known locations across 16 categories in 184 countries worldwide. Unlike most existing remote sensing vision-language datasets that only provide basic descriptions of current scenes, our RSWK dataset includes richer remote sensing domain expertise and world knowledge about the objects within these scenes. For instance, from a remote sensing perspective, it provides information on surface reflectance, spectral indices, and atmospheric conditions. From a world knowledge perspective, it includes historical background, cultural significance, construction period, and major events. This combination of remote sensing expertise and world knowledge not only enhances the RSWK dataset’s utility for visual analysis of remote sensing images, but also provides the model with deeper semantic context, overcoming the limitations of traditional datasets and enabling remote sensing VLMs to perform more complex cognitive tasks. Furthermore, our dataset incorporates historical, cultural, and social backgrounds from various countries and regions, allowing VLMs to be trained across diverse geographical and cultural contexts, thereby improving their generalization ability and understanding of different cultural settings.
@@
Figure 1: The construction process of the Remote Sensing World Knowledge (RSWK) dataset begins with collecting landmark data from around the world, followed by extracting geographic information to pinpoint precise location coordinates. Using these coordinates, remote sensing images are acquired, which are further standardized through image processing techniques. Corresponding remote sensing expert knowledge, such as surface temperature, climate, atmospheric conditions, and spectral indices, is also included. Additionally, world knowledge is retrieved from online resources, providing detailed background information about the landmarks, including historical context, cultural significance, and major events. This combined information is structured into organized attributes to align image and text data, forming a final multimodal dataset. The resulting RSWK dataset integrates high-resolution images with extensive remote sensing and world knowledge, enabling advanced semantic understanding in remote sensing applications.
@@
Furthermore, to effectively leverage the comprehensive information provided by the RSWK dataset, we propose the Remote Sensing Retrieval-Augmented Generation (RS-RAG) model. RS-RAG is designed to enhance the capacity of VLMs to generate contextually enriched and knowledge-grounded responses for remote sensing imagery. It operates by integrating external knowledge, both domain-specific and general world knowledge, retrieved from a multimodal knowledge base constructed using the RSWK dataset. The model consists of two main components: (1) the Multi-Modal Knowledge Vector Database Construction module, which encodes satellite imagery and textual descriptions into a shared embedding space using unified image and text encoders to enable efficient cross-modal retrieval; and (2) the Knowledge Retrieval and Response Generation module, which retrieves top-ranked knowledge entries based on image and text queries, re-ranks them via a fused similarity score, and incorporates the selected content into a knowledge-augmented prompt. This prompt is then passed to the vision-language model, enabling it to generate responses that go beyond semantic understanding and reflect deeper background knowledge. By coupling visual input with relevant contextual information, RS-RAG significantly improves the interpretability and accuracy of vision-language outputs, particularly for complex queries involving geospatial, historical, or environmental reasoning.
To assess the effectiveness of the proposed model, we construct a comprehensive benchmark and evaluate performance across three representative tasks: image captioning, image classification, and visual question answering (VQA). To better leverage the rich semantics of the dataset, we further refine the VQA task into three distinct subtypes that target: (1) comprehensive reasoning, (2) domain-specific remote sensing understanding, and (3) general world knowledge interpretation. Results on these tasks demonstrate the model’s ability to produce accurate, context-rich descriptions, deliver semantically informed scene classifications, and provide precise answers to knowledge-intensive queries by leveraging both visual and textual modalities. Through these findings, RS-RAG demonstrates its potential to substantially advance the capabilities of VLMs in remote sensing, effectively bridging the gap between imagery and comprehensive contextual knowledge. Our main contributions are summarized as follows:
@@
  • We construct the Remote Sensing World Knowledge (RSWK) dataset, a large-scale multimodal dataset containing 14,820 high-resolution remote sensing images and rich textual descriptions of globally recognized landmarks across 16 categories from 184 countries. The descriptions incorporate both domain-specific remote sensing knowledge and general world knowledge.
@@
  • We design a benchmark spanning image captioning, classification, and VQA, with VQA further divided into three subtypes to assess comprehensive reasoning, remote sensing expertise, and world knowledge. This setup facilitates systematic evaluation of both semantic perception and knowledge-intensive reasoning.
  • Extensive experiments demonstrate that RS-RAG consistently outperforms state-of-the-art VLMs across all tasks, particularly on queries requiring external world knowledge. These results highlight the effectiveness of RS-RAG in bridging remote sensing imagery with structured knowledge, and point to promising future directions for research on remote sensing vision-language models.
@@ II-A Remote Sensing Multimodal Datasets
Several multimodal datasets [24, 41, 30, 2, 35] have been developed to bridge the gap between vision and language in the remote sensing domain. UCM Captions [29], Sydney Captions [29], and RSICD [26] are among the earliest datasets that provide textual descriptions for remote sensing images. Each image in these datasets is paired with five relatively simple human-written sentences, offering only a basic level of semantic information. To enhance the quality and richness of textual annotations, RSGPT [10] recently introduced RSICap, a high-quality image captioning dataset with detailed human-annotated descriptions of aerial scenes. RSICap serves as a valuable resource for fine-tuning and developing domain-specific vision-language models in remote sensing. Beyond manual annotation, researchers have explored the construction of large-scale datasets using automatic or hybrid approaches applied to existing data sources. For instance, RS5M [39] was created by aggregating 11 publicly available image–text paired datasets along with three large-scale class-level labeled datasets. Captions were generated using BLIP-2, resulting in a diverse and large-scale dataset suitable for training foundational multimodal models. Similarly, RemoteCLIP [22] compiled a large-scale dataset by integrating 10 object detection datasets, 4 semantic segmentation datasets, and 3 remote sensing image–text datasets, enabling contrastive pretraining for cross-modal alignment. In addition, GeoChat [14] introduced the RS Multimodal Instruction Dataset, which incorporates heterogeneous data sources, including three object detection datasets, one scene classification dataset, and a visual question answering dataset focused on flood detection. Recent datasets such as SkyScript [32], LhrsAlign [27], and ChatEarthNet [34] leverage external geospatial sources, such as OpenStreetMap, along with geo-coordinates to construct large-scale, globally distributed multimodal benchmarks. More recently, FedRSCLIP [20] proposed a new multimodal remote sensing dataset specifically designed for federated learning scenarios, further expanding the applicability of vision-language research in distributed settings. These efforts collectively advance the development of remote sensing VLMs by providing diverse, rich, and large-scale multimodal data resources tailored to various downstream tasks.
@@ II-B Remote Sensing Multimodal Model
With the growing availability of remote sensing multimodal datasets, a number of vision-language models [10, 4, 28, 37, 15, 36] have been proposed to enhance the understanding and interpretation of aerial imagery through natural language. [16] provides the first comprehensive review of vision-language models in remote sensing, systematically summarizing tasks, datasets, and methods, and identifying key challenges and future directions. RSCLIP [17] is a pioneering vision-language model designed for remote sensing scene classification, which leverages contrastive vision-language supervision and incorporates a pseudo-labeling technique along with a curriculum learning strategy to improve zero-shot classification performance through multi-stage fine-tuning. RSGPT [10] represents one of the earliest attempts to build a generative pretrained model for remote sensing. It fine-tunes only the Q-Former and a linear projection layer on the proposed RSICap dataset, achieving notable improvements in both image captioning and visual question answering tasks. Similarly, GeoChat [14] adapts the LLaVA-1.5 architecture and fine-tunes it on its proposed remote sensing multimodal instruction-following dataset, offering multi-task conversational capabilities grounded in high-resolution satellite imagery. In addition, to address the common issue of hallucination in remote sensing vision-language models, a Helpful and Honest Remote Sensing Vision-Language Model [28], named H2RSVLM, is proposed and fine-tuned on the RSSA dataset, the first dataset specifically designed to enhance self-awareness in remote sensing VLMs. More recently, RS-MoE [18] was proposed as the first Mixture-of-Experts-based VLM tailored for remote sensing. It features a novel instruction router that dynamically dispatches tasks to multiple lightweight expert LLMs, allowing each expert to specialize in a specific subset of tasks.
Figure 2: Overview of the RSWK dataset. (a) Global distribution of landmarks used in the dataset, with color indicating the number of landmarks per country. (b) Statistical summaries of landmark counts across the top 50 countries (left) and all 16 categories (right). (c) A specific example from the RSWK dataset, showcasing the Sydney Opera House, including its satellite imagery, remote sensing domain knowledge, and structured world knowledge.
II-C Retrieval-Augmented Generationl
To address the limitations of VLMs in handling knowledge-intensive or domain-specific tasks, Retrieval-Augmented Generation (RAG) [42] has emerged as a promising solution for enhancing the accuracy and reliability of generative models by incorporating information from external, relevant knowledge sources. By retrieving and integrating contextual knowledge during inference, RAG significantly improves the reasoning ability and factual consistency of large models. In recent years, researchers have explored the integration of RAG into a variety of domains, including natural language processing[7], natural image understanding[25], and video analysis [12]. However, the application of RAG in the remote sensing domain remains limited. More recently, RAGCap [3] proposes a fine-tuning-free method for remote sensing image captioning, where relevant image-caption pairs are retrieved from a training set to guide generation. Similarly, ImageRAG [38] introduces an image-contextual retrieval mechanism based on RAG principles, which transforms ultra-high-resolution remote sensing image analysis into a long-context selection task, enabling efficient understanding with minimal training data.
@@ III Dataset Construction
Most existing remote sensing multimodal datasets focus primarily on basic descriptions of objects in a scene, typically providing only information about the objects present. While these datasets serve a purpose in supporting fundamental tasks such as scene classification and image captioning, they fall short in applications requiring complex semantic understanding or contextual awareness. To address this limitation, in our paper, we propose the Remote Sensing World Knowledge (RSWK) dataset, which encompasses high-resolution remote sensing imagery of well-known locations worldwide, along with domain knowledge and world knowledge described in natural language. This dataset not only fills the gap in knowledge depth and breadth found in current remote sensing datasets but also provides a new foundation for advancing remote sensing technology toward intelligent applications. The core value of remote sensing data lies in its ability to deliver rich geographic and spatial distribution information. By integrating high-resolution imagery from across the globe with detailed domain knowledge and world knowledge, the RSWK dataset extends this value, expanding the application of remote sensing data from traditional foundational tasks to complex scenarios that require deeper semantic understanding.
III-A Data Collection and Processing
Fig. 1 illustrates the end-to-end pipeline for constructing the RSWK dataset, which aims to bridge the gap between remote sensing domain knowledge and encyclopedic world knowledge of globally distributed landmarks. The pipeline comprises four key stages that systematically curate and align multimodal information from global sources, integrating high-resolution remote sensing imagery, remote sensing expert knowledge, and contextual world knowledge.
1) Worldwide Landmark Data Collection: The first stage involves generating a comprehensive list of globally recognized landmarks using GPT-4o [11]. To ensure broad coverage across geographic regions and cultural contexts, we prompt the model to identify representative landmarks spanning a wide range of categories, including transportation hubs (e.g., airports), recreational areas (e.g., amusement parks, beaches), cultural institutions (e.g., museums, theaters, universities), historical sites, and natural environments (e.g., forests).
2) Remote Sensing Imagery Acquisition: The second stage focuses on acquiring high-resolution satellite imagery for each landmark. We first utilize the Google Geocoding Maps API to obtain precise geographic coordinates and region-level viewpoint information for each target location. To resolve potential ambiguities in place names, we implement a cross-source validation mechanism that references multiple geospatial data providers and applies regional constraints to enhance geolocation accuracy. For image acquisition, we access the ArcGIS World Imagery Tile Map Service provided by Esri, which delivers high-resolution satellite imagery in standard XYZ tile format using the Web Mercator projection. Based on the retrieved geo-coordinates, we collect imagery tiles at appropriate zoom levels ranging from 16 to 18, depending on the spatial extent of each landmark. This adaptive zoom-level strategy ensures that landmarks of varying scales (e.g., airport, forest) are fully captured within the resulting imagery. To enable efficient downloading, we implement a multithreaded queue-based retrieval system that parallelizes tile requests across locations. The downloaded tiles are then mosaicked into seamless, high-resolution scene-level images. Finally, we apply standardized normalization, cropping, and resizing operations to ensure consistent image quality, resolution, and format across the dataset.
3) Remote Sensing Knowledge Gathering and Structuring: The third stage focuses on acquiring structured remote sensing expert knowledge using Google Earth Engine (GEE)[40]. This knowledge spans a broad set of satellite-derived physical and environmental variables. These include surface reflectance bands, vegetation and land surface indices, as well as biophysical parameters like surface albedo, emissivity, and land cover classifications. In addition, the dataset incorporates a range of meteorological and geophysical variables obtained from authoritative products such as MERRA-2, HLS, and ERA5. These include land surface temperature, snowmelt flux, total precipitation, surface wind speed and direction, solar angles, and pressure-related measurements. To ensure temporal consistency between imagery and knowledge fields, all dynamic variables are aggregated either annually or monthly within the same year as the image acquisition.
4) World Knowledge Retrieval: The final stage involves curating structured world knowledge to provide rich semantic context for each landmark. We leverage the Wikipedia API to extract descriptive entries for each site, encompassing diverse dimensions such as historical background, architectural characteristics, cultural significance, and major events. To ensure the relevance and quality of the extracted content, we apply a multi-step post-processing pipeline. Specifically, keyword-based filtering and entity recognition are used to discard generic or off-topic segments, while DeepSeek [21] is employed to refine textual content by correcting formatting inconsistencies, improving coherence, and removing noisy or redundant information. The resulting knowledge is clean, semantically rich, and structured into predefined fields, enabling effective integration with downstream vision-language tasks that require contextual understanding and factual grounding.